Circular No. 9078

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2009 QG_{31}

An apparently asteroidal object discovered on CCD images taken remotely in the course of the “La Sagra Sky Survey” (LSSS), with a 0.45-m f/2.8 reflector located at Sagra mountain in southeastern Spain, has been found to show cometary appearance by CCD astrometrists elsewhere. The observation tabulated below was assigned “discovery” status after the Minor Planet Center linked “one-night” LSSS observations made on Aug. 19 with others made on Aug. 25 (though earlier LSSS observations were made on Aug. 16), and later LSSS and Catalina observations linked by the MPC allowed the issuance a comet-like orbit on Sept. 8 (MPEC 2009-R26).

2009 UT  α_{2000}  δ_{2000}  Mag.
Aug. 19.04841  22^{h}45^{m}37.^{s}88  −10^{o}10′38′′.0  18.4

Images obtained with a 0.40-m reflector on Sept. 10.0 UT by the LSSS team, to look for cometary appearance, showed no obvious coma, but images taken by A. F. Tubbiolo and R. S. McMillan with the Spacewatch 1.8m reflector on Sept. 12.3 show the object to be diffuse with a coma ≈ 12′′–15′′ in diameter. E. Bryssinck (near Cloudcroft, NM, U.S.A.; Tzec Maun 0.4-m reflector; Sept. 16.3) finds a coma diameter of ∼ 10′′. F. Hormuth, Almería, Spain, writes that three stacked 180-s images obtained with the Calar Alto 3.5-m telescope (+ prime-focus camera ‘LAICA’ + r′ filter) on Sept. 17.9 show a nearly round coma of diameter 12′′. R. Miles, Stourton Caundle, Dorset, England, reports that images taken on Sept. 18.4 by G. Muler with the 2-m ‘Faulkes Telescope North’ at Haleakala show a faint coma of diameter ∼ 10′′. G. Sostero, E. Guido, P. Camilleri, M. Jaeger, W. Vollmann, and E. Prosperi report that 24 stacked, unfiltered, 60-s exposures obtained remotely with a 0.35-m reflector at the Skylive-Grove Creek Observatory (near Trunkey, Australia) on Sept. 18.5 show a round coma nearly 7′′ in diameter. R. Ligustri (Udine, Italy, 0.35-m reflector; Sept. 19.9) measured a coma diameter of 10′′–15′′ from fifteen stacked 240-s images. Elliptical orbital elements from MPEC 2009-S67 from 53 observations, Aug. 16–Sept. 19 (the available astrometry appears on MPS 292728, 294488, 295077, and 295950):

T = 2009 Oct. 10.4284 TT  ω = 5.^{o}9231
e = 0.399346  Ω = 346.2506 }2000.0
q = 2.149305 AU  i = 5.0628
a = 3.578274 AU n^{o} = 0.1456108 P = 6.77 years
2009 September 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green